OpenStudy (anonymous):

what is the slope of the line represented by the equation y= 1/2x-4?

OpenStudy (wasiqss):

1/2

OpenStudy (wasiqss):

n u havnt used parenthisis correctly, u shud have written y=(x-2)/2

OpenStudy (anonymous):

(1/2) is the slope for y = (1/2)x-4